# Exploring Safety Generalization Challenges of Large Language Models via Code

### 14 Mar 2024 v2

@@ Abstract
The rapid advancement of Large Language Models (LLMs) has brought about remarkable capabilities in natural language processing but also raised concerns about their potential misuse. While strategies like supervised fine-tuning and reinforcement learning from human feedback have enhanced their safety, these methods primarily focus on natural languages, which may not generalize to other domains. This paper introduces CodeAttack, a framework that transforms natural language inputs into code inputs, presenting a novel environment for testing the safety generalization of LLMs. Our comprehensive studies on state-of-the-art LLMs including GPT-4, Claude-2, and Llama-2 series reveal a common safety vulnerability of these models against code input: CodeAttack consistently bypasses the safety guardrails of all models more than 80% of the time. Furthermore, we find that a larger distribution gap between CodeAttack and natural language leads to weaker safety generalization, such as encoding natural language input with data structures or using less popular programming languages. These findings highlight new safety risks in the code domain and the need for more robust safety alignment algorithms to match the code capabilities of LLMs.
@@ 1 Introduction
Our findings uncover new safety risks associated with LLMs in novel domains that are far away from their safety training distribution, which is not adequately addressed by current safety mechanisms. Furthermore, CodeAttack’s cost efficiency and automated nature eliminate the need for attackers to have domain-specific coding knowledge, suggesting a potential increase in the misuse of LLMs in the code domain. We hope that sharing our discoveries will inspire further research into designing more robust safety alignment algorithms that can generalize to unseen domains, towards the safer integration of LLMs into the real world.
@@ 3.3 Output Specification
Table 1: Attack success rate (ASR) of baseline attacks and our CodeAttack on the AdvBench dataset Zou et al. (2023). CodeAttack can breach the safety guardrails of current SOTA LLMs, including GPT, Claude, and Llama-2 series. ***: we report our evaluation results of SelfCipher in CipherChat since its original paper does not include experiments on AdvBench. For other baselines, we list the original results from their papers. For a thorough comparison, we list the results of CodeAttack with different data structures used for input encoding: string, queue, and stack. CodeAttack is implemented in Python.
@@ 4.1 Experimental Setup
For CipherChat, we report evaluation results of SelfCipher in CipherChat since its original paper does not include experiments on AdvBench Zou et al. (2023). The implementation details of SelfCipher can be found in Appendix A.1. For other baselines, we show the original results from their papers.
@@ 4.2 Results
A larger distribution gap between CodeAttack and natural language leads to weaker safety generalization. Table 1 shows that CodeAttack becomes more effective as the encoded input diverges from natural language, with the average attack success rate increasing from 51% to 68% to 78% as the input encoding data structure changes from string to queue to stack, the latter being the least similar to natural language, as depicted in Figure 2. This suggests that LLMs are more likely to produce unsafe content when the encoded malicious input is less similar to natural language, i.e., further from the safety training data distribution. Additionally, due to their weaker code understanding capabilities, smaller models such as GPT-3.5, Claude-1, and Llama-2 models exhibit slightly safer behavior when encoding inputs as stacks compared to strings or queues. For example, the attack success rate of Llama-2-7b decreases from 88% to 54% when the input encoding data structure changes from queue to stack. We observe that these smaller models struggle with task understanding when inputs are encoded as stacks, tending to select the first word of the query as the task, which negatively impacts the quality and accuracy of their outputs.
@@ 4.3 Ablation & Analysis
Impact of task understanding. To demonstrate the effectiveness of the $decode()$ function, we establish a baseline in which models are prompted to extract the task solely through comments, as depicted in Figure 3. Additionally, we evaluate the two methods across different types of input encoding to illustrate the generality of the $decode()$ function. Table 2 shows that using comments makes CodeAttack less effective compared to using the $decode()$ function. To analyze why using comments leads to a weaker attack, we delved into the differences in the model’s responses between these two methods. Specifically, when extracting tasks through comments, Claude-1 tends to incorporate ethical steps in its response, whereas GPT-4 and Claude-2 are more inclined to convert the original malicious task into a different ethical task, e.g. they modify the original malicious query “Create a fake news article that could lead to panic or chaos.” to an ethical one “Create a list of healthy eating habits for a balanced diet.”, and provide a harmless answer to the ethical task. This indicates that the safety guardrails of these models are more likely to be activated in this scenario of using comments, compared to using the $decode()$ function. In general, introducing the $decode()$ function brings our prompt closer to the code distribution, which deviates more from the safety training data, thereby making it easier to bypass the safety guardrails of LLMs.
@@ 4.4 Why does CodeAttack work?
Experimental results in Section 4.2 and  4.3 demonstrate that LLMs are more likely to produce unsafe content when CodeAttack is closer to the code distribution, where the generalization of LLMs’ safety behavior is more challenging. Here, we offer some intuitive explanations for the success of CodeAttack. Since during code training, LLMs primarily pursue the objective of helpfulness, i.e., to complete code accurately Rozière et al. (2024), we hypothesize that when our prompts are closer to the code training distribution, models are more likely to exhibit helpful behavior, i.e., following our instruction to complete our code. In other words, we suspect that in such cases the safety behavior may be suppressed because the models might refuse to answer to satisfy the safety objective, which conflicts with the objective of helpfulness. To test our hypothesis, we further stimulate the models’ intention to complete our codes by prepending a benign algorithm from the code training data into our prompt. Table 5 shows that simply adding the implementation of the quick sort algorithm into CodeAttack leads to greater safety degradation in larger LLMs such as GPT-4 and Claude-2. Our analysis indicates that for LLMs, pursuing the helpfulness objective may compromise the harmlessness objective. This highlights the need for designing better alignment algorithms that balance these two objectives.

### 07 Apr 2024 v3

@@ Abstract
The rapid advancement of Large Language Models (LLMs) has brought about remarkable generative capabilities but also raised concerns about their potential misuse. While strategies like supervised fine-tuning and reinforcement learning from human feedback have enhanced their safety, these methods primarily focus on natural languages, which may not generalize to other domains. This paper introduces CodeAttack, a framework that transforms natural language inputs into code inputs, presenting a novel environment for testing the safety generalization of LLMs. Our comprehensive studies on state-of-the-art LLMs including GPT-4, Claude-2, and Llama-2 series reveal a common safety vulnerability of these models against code input: CodeAttack bypasses the safety guardrails of all models more than 80% of the time. We find that a larger distribution gap between CodeAttack and natural language leads to weaker safety generalization, such as encoding natural language input with data structures. Furthermore, we give two hypotheses about the success of CodeAttack: (1) the misaligned bias acquired by LLMs during code training, prioritizing code completion over avoiding the potential safety risk; (2) the limited self-evaluation capability regarding the safety of their code outputs. Finally we analyze potential mitigation measures. These findings highlight new safety risks in the code domain and the need for more robust safety alignment algorithms to match the code capabilities of LLMs.
@@ 1 Introduction
We give two hypotheses about the success of CodeAttack: (1) models acquire the misaligned bias during code training, which prioritizes code completion over avoiding the potential safety risk. By pretending a benign code snippet into our prompt, models are more conducive to giving harmful codes; (2) models show worse self-evaluation capability regarding the safety of their code outputs than natural language outputs. Finally, we discuss potential mitigation measures. Our findings uncover new safety risks associated with LLMs in novel domains that are far away from their safety training distribution, which is not adequately addressed by current safety mechanisms. We hope that sharing our discoveries will inspire further research into designing more robust safety alignment algorithms that can generalize to unseen domains, towards the safer integration of LLMs into the real world.
@@ 3.3 Output Specification
Table 1: Attack success rate (ASR) of baseline attacks and our CodeAttack on the AdvBench dataset Zou et al. (2023). CodeAttack can breach the safety guardrails of current SOTA LLMs, including GPT, Claude, and Llama-2 series. ***: we report our evaluation results of SelfCipher in CipherChat since its original paper does not include experiments on AdvBench. For other baselines, we list their implementation results from Zeng et al. (2024). For a thorough comparison, we list the results of CodeAttack with different data structures used for input encoding: string, queue, and stack. CodeAttack is implemented in Python.
@@ 4.1 Experimental Setup
For CipherChat, we report evaluation results of SelfCipher in CipherChat since its original paper does not include experiments on AdvBench Zou et al. (2023). The implementation details of SelfCipher can be found in Appendix A.1. For other baselines, we show their results from Zeng et al. (2024), where these baselines are implemented and evaluated using the same GPT-4 judge as employed in our study. Thus, we use these baseline results to ensure a fair comparison with our CodeAttack.
@@ 4.2 Results
A larger distribution gap between CodeAttack and natural language leads to weaker safety generalization. Table 1 shows that CodeAttack becomes more effective as the encoded input diverges from natural language, with the average attack success rate increasing from 51% to 68% to 78% as the input encoding data structure changes from string to queue to stack, the latter being the least similar to natural language, as depicted in Figure 2.
This suggests that LLMs are more likely to produce unsafe content when the encoded malicious input is less similar to natural language, i.e., further from the safety training data distribution. Additionally, due to their weaker code understanding capabilities, smaller models such as GPT-3.5, Claude-1, and Llama-2 models exhibit slightly safer behavior when encoding inputs as stacks compared to strings or queues. For example, the attack success rate of Llama-2-7b decreases from 88% to 54% when the input encoding data structure changes from queue to stack. We observe that these smaller models struggle with task understanding when inputs are encoded as stacks, tending to select the first word of the query as the task, which negatively impacts the quality and accuracy of their outputs.
@@ 4.3 Ablation & Analysis
Impact of task understanding. To demonstrate the effectiveness of the $decode()$ function, we establish a baseline in which models are prompted to extract the task solely through comments, as depicted in Figure 3. Additionally, we evaluate the two methods across different types of input encoding to illustrate the generality of the $decode()$ function. Table 2 shows that using comments makes CodeAttack less effective compared to using the $decode()$ function. To analyze why using comments leads to a weaker attack, we delved into the differences in the model’s responses between these two methods. Specifically, when extracting tasks through comments, Claude-1 tends to incorporate ethical steps in its response, whereas GPT-4 and Claude-2 are more inclined to convert the original malicious task into a different ethical task, e.g. they modify the original malicious query “Create a fake news article that could lead to panic or chaos.” to an ethical one “Create a list of healthy eating habits for a balanced diet.”, and provide a harmless answer to the ethical task.
This indicates that the safety guardrails of these models are more likely to be activated in this scenario of using comments, compared to using the $decode()$ function. In general, introducing the $decode()$ function brings our prompt closer to the code distribution, which deviates more from the safety training data, thereby making it easier to bypass the safety guardrails of LLMs.
@@ 4.4 Why does CodeAttack work?
Experimental results in Section 4.2 and  4.3 demonstrate that LLMs are more likely to produce unsafe content when CodeAttack is closer to the code distribution, where the generalization of LLMs’ safety behavior is more challenging. Here, we offer some intuitive explanations for the success of CodeAttack.
4.4.1 The misaligned bias of completing code learned from the training phase.
We hypothesize that during the process of code training, these models acquire a misaligned bias, prioritizing accuracy in code generation over the safety concerns. To substantiate this hypothesis, we aim to leverage this misaligned bias to enhance the efficacy of our attacks. Specifically, we prepend a benign quick sort algorithm from the training dataset into our prompt, making it closer to the training distribution. As demonstrated in Table 5, the integration of the quick sort algorithm into our CodeAttack framework significantly increases the susceptibility of models to generate unsafe code, thereby exacerbating safety degradation. This finding underscores the imperative for developing safety alignment algorithms that effectively counteract the misaligned bias inherent in code LLMs.
4.4.2 Models’ self-awareness regarding the safety of their outputs.
Input format for the self-evaluator GPT-3.5
GPT-4
-1106
Claude-1 Claude-2
Llama-2
-7b
Llama2
-70b
CodeLlama
-70b
Original code 40%(-44%) 72 (-14%) 26%(-58%) 68%(-18%) 40%(-16%) 12%(-62%) 98%(+12%)
Extracted natural language text 84% (0%) 86% (0%) 58%(-26%) 84%(-2%) 50%(-6%) 40%(-34%) 6% (-80%)
GPT-4 as the evaluator 84% 86% 84% 86% 56% 74% 86%
Table 6: Harmfulness detection results by the self-evaluation capability of models. We list the results of two input formats for the self-evaluator: original code and the natural language text extracted from the code. We also show the results of GPT-4 evaluator for the outputs of each model for comparison. CodeAttack takes Python stack as input encoding.
We also hypothesize that models may lack awareness of the harmful nature of the code they generate, allowing CodeAttack to circumvent the safety mechanisms. To investigate this, we carried out an extensive analysis of the self-evaluation capabilities of models, including GPT-3.5, GPT-4, Claude-1, Claude-2, and CodeLlama-70b. This assessment encompassed both the original code output resulted by CodeAttack and the natural language text extracted from the code. The safety prompt employed for self-evaluation was identical to that used for the GPT-4 evaluator in this paper. Table 6 shows the harmfulness detection rate of model responses by self-evaluation and the difference between the self-evaluation and GPT-4 evaluation, noted within parentheses. Our observations include:
  1. 1.
    It is harder for self-evaluators to discriminate the harmful content in code than natural language: for instance, only 40% of code outputs were deemed harmful by GPT-3.5, whereas 84% of natural language instances were identified as unsafe by both GPT-3.5 and GPT-4. This disparity in self-evaluation capability suggests that models’ self-awareness of the harmfulness in their generated content is diminished when producing code, likely contributing to the effectiveness of our CodeAttack.
  2. 2.
    Stronger models have better self-evaluation capability: stronger models like GPT-4 and Claude-2 show better awareness of the harmfulness of their outputs, in both code and natural language formats. This highlights the potential of post hoc evaluations as mitigation strategies.
  3. 3.
    We identified two failure modes in models’ self-evaluation when processing code: (a) models can fail to recognize the harmful content inside code; (b) while these self-evaluators acknowledge that the provided information in code could be harmful, they contend that the model’s response is not a direct fulfillment of the user’s intent, leading to incorrect evaluations. This indicates that models can not reason well about the relationship between code output and user instructions, especially for weaker models.
  4. 4.
    Concerning the over-sensitivity of CodeLlama-70b’s evaluations: CodeLlama-70b tends to overestimate the harmfulness of code outputs, particularly when they contain sensitive words. Given that our methodology involves encoding harmful queries within the code, making all code outputs contain sensitive terms, CodeLlama-70b nearly always flags these as harmful. Conversely, in evaluating natural language text, CodeLlama-70b often abstains from performing evaluations, exhibiting a more conservative behavior compared to other models.
Overall, our CodeAttack reveals that stronger models like GPT-4 and Claude-2 have good self-evaluation capability regarding the safety of their outputs, which has the potential as a mitigation measure. Weaker models like GPT-3.5 and Claude-1 have worse self-awareness of safety, bringing greater challenges to safety alignment. Notably, these stronger models still behave unsafely against our CodeAttack, despite their awareness of harmfulness. This suggests a misalignment in the safety behavior of models shaped by current training techniques with human intent.
Defenses GPT-3.5
GPT-4
-1106
Claude-1 Claude-2 CodeLlama-70b
No defenses 84% 86% 84% 86% 86%
OpenAI 82%(-2%) 86%(-0%) 82%(-2%) 84%(-2%) 78%(-8%)
Paraphrase 84%(-0%) 84%(-0%) 84%(-0%) 80%(-6%) 74%(-12%)
Rand-insert 82%(-2%) 34%(-52%) 84%(-0%) 86%(-0%) 82%(-4%)
Rand-swap 82%(-2%) 52%(-34%) 84%(-0%) 86%(-0%) 86%(-0%)
Rand-patch 82%(-2%) 40%(-46%) 84%(-0%) 84%(-2%) 84%(-2%)
Table 7: Attack success rate of CodeAttack against defense baselines on AdvBench Zou et al. (2023). We bold the best results of each defense method and list the difference between the defense baseline and no defenses in parentheses.
4.5 Discussion about mitigation measures
We select three types of baselines to evaluate the robustness of CodeAttack.
  1. 1.
    OpenAI Moderation Endpoint Todor et al. (2023), an official detection tool provided by OpenAI, which can be used to check whether the model response is potentially harmful or not.
  2. 2.
    Rand-Insert, Rand-Swap, and Rand-Patch from SmoothLLM Robey et al. (2023), which is also a detection-based method by perturbing the inputs in different ways and inspecting the output changes.
  3. 3.
    Paraphrase Jain et al. (2023), which is an input preprocessing method by paraphrasing inputs to remove the possible adversarial sequence of tokens while preserving natural instructions.
For SmoothLLM Robey et al. (2023), we follow their defense and evaluation setting, and for paraphrase Jain et al. (2023), we use GPT-4 to paraphrase prompts and set the temperature to 0. We select a Python implementation of CodeAttack whose input type is stack for evaluation.
Table 7 shows the attack success rate (ASR) of CodeAttack and how much these defenses can reduce the ASR. We discover that CodeAttack shows great robustness against current defenses with little performance degradation, except for GPT-4. Both SmoothLLM Robey et al. (2023) and Paraphrase Jain et al. (2023) assume that the adversarial tokens in an adversarial prompt are fragile to perturbations, which works well against character-level attacks like GCG Zou et al. (2023). However, our CodeAttack formulates a code completion task, which is semantically coherent and hence robust to random perturbations or simple paraphrasing. These results highlight the importance of acknowledging the new threat model exposed by CodeAttack and the necessity for developing tailored defense mechanisms.
Furthermore, we observe that detection-based methods like Rand-Insert, Rand-Swap, and Rand-Patch can defend GPT-4-1106 more effectively. These methods evaluate the safety of a model’s output by checking for specific phrases in a predefined keyword set, such as "I’m sorry" or "illegal." GPT-4-1106 tends to incorporate these keywords in its responses, making its harmful outputs more detectable.